If I need $4000 in four years and start saving $1000 today what rate of return would I need? 4000 = 1000(1 plus X)4 I tried 4000 plus 1 4001/1001 and got 4 I know the answer should be around 41 but not sure how to slove this equation.
Answer by mananth(16949)

i= (fv/pv)^(1/n) -1

fv= 4000
pv= 1000
n= 4
i=(4000/1000)^1/n-1
i= 4^1/4-1
i=0.41
i=41 %